### v1

# Adaptive Trajectory Bundle Method for Roll-to-Roll Manufacturing Systems

Abstract
Roll-to-roll (R2R) manufacturing demands precise tension and velocity control under strict operational constraints. Model predictive control requires gradient computation, while sampling-based methods such as MPPI struggle with hard constraint satisfaction. This paper presents an adaptive trajectory bundle method that achieves rigorous constraint handling through derivative-free sequential convex programming. The approach approximates nonlinear dynamics and costs via interpolated sample bundles, with adaptive trust regions and penalty parameters ensuring robust convergence without manual tuning. Simulations on a six-zone R2R system demonstrate tracking accuracy comparable to gradient-based MPC with superior constraint satisfaction over sampling-based alternatives.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
Roll-to-roll (R2R) manufacturing processes are essential for high-throughput production of flexible electronics, photovoltaics, functional films, and advanced materials. These systems transport continuous web material through sequential processing zones, where precise tension and velocity regulation are critical to product quality. Excessive tension causes web breakage, while insufficient tension leads to wrinkling and defects. The control challenge is compounded by strong coupling between adjacent zones, time-varying parameters due to changing roll radii, and strict operational constraints on tensions, velocities, and actuator rates.
Traditional control approaches face fundamental limitations. Model Predictive Control (MPC), while capable of handling constraints, requires accurate gradient information and suffers from computational burden in high-dimensional settings. Sampling-based methods like Model Predictive Path Integral (MPPI) control avoid gradient computation but struggle with hard constraint satisfaction. This motivates the need for a framework combining constraint handling with derivative-free optimization.
I-A Related Work
Classical R2R control relies on PID, LQR, and MPC variants to regulate web tension and transport speed. Foundational modeling established tension/velocity decoupling strategies [koc2002winding], while industrial studies refined adaptive PI/PID implementations for operating-point drift [raul2015adaptivePI, chen2017fuzzy]. Recent work emphasizes advanced control methods [martin2024stabilization, martin2025sequential] and physics-consistent tension models [jeong2021tension, he2024multispan]. Manufacturing surveys underline the role of monitoring and feedback design in achieving system robustness [martin2022hinfty, martin2025hinfty]. Data-driven approaches report robust constrained tension regulation [chen2023robustR2R] and AI-assisted tension estimation [gafurov2024webtension, gafurov2025aidt], with reviews pointing to hybrid controllers as a promising direction [martin2025review].
Despite these advances, classical methods require extensive tuning and provide limited guarantees under coupled nonlinear dynamics. Advanced MPC formulations demand gradient computation and careful initialization, with computational requirements becoming prohibitive for systems with many zones or long prediction horizons.
MPPI control is inherently unconstrained, typically relying on soft-constraint penalties [park2025csc, yin2023shield]. Methods to enforce hard constraints include CBF-based shields [yin2023shield], though these can be myopic [rabiee2024guaranteed]. Safe-by-construction approaches such as GS-MPPI embed composite CBFs into dynamics [rabiee2024guaranteed], while MPPI-DBaS augments states with discrete barrier states [wang2025mppi]. DualGuard-MPPI integrates Hamilton-Jacobi reachability for provable safety [borquez2025dualguard]. Other approaches improve sample efficiency using learned surrogates (BC-MPPI) [ezeji2025bc], specialized samplers [yan2023output, mpopi_legged_2025], or clustering algorithms to address multimodality [park2025csc].
While these MPPI variants succeed in robotics, they face limitations for R2R manufacturing. The single-shooting formulation makes long-horizon planning unstable for coupled multi-zone dynamics, and existing constraint mechanisms either provide soft guarantees, require additional filtering layers, or demand specialized model structures.
Recently, the Trajectory Bundle Method (TBM) introduced a derivative-free optimal control paradigm based on sequential convex programming [tracy2025trajectory]. Instead of Taylor-series approximations, TBM uses interpolated bundles of sampled trajectories to approximate cost, dynamics, and constraint functions.
I-B Contributions
This paper develops an adaptive trajectory bundle method for constrained R2R control. The main contributions are:
1) We derive the complete mathematical formulation applying TBM to R2R tension-velocity dynamics, including physics-based cost functions, operational constraints, and asymmetric risk handling for over/under-tensioning. The overall workflow of Algorithm 1 is illustrated in Figure 1.
2) We introduce adaptive trust region management and penalty scheduling that automatically adjust based on constraint violation metrics, reducing the need for manual tuning.
3) We establish convergence guarantees for the adaptive TBM, proving that the algorithm achieves feasibility in finite iterations and converges to stationary points of the constrained R2R control problem.
The remainder of this paper is organized as follows. Section II derives R2R system dynamics. Section III presents the adaptive TBM theoretical framework. Section IV establishes convergence guarantees. Section V provides simulation results. Section VI concludes the paper.
Initialize $z^{0},\Delta^{0},\mu^{0},\boldsymbol{\gamma}^{0}$Bundle GenerationBuild Bundle MatricesSolve Convex SubproblemRecover TrajectoryCompute ViolationsConverged?Return $z^{*}$Trust Region AdaptPenalty AdaptYesNo
Figure 1: Flowchart of the Adaptive Trajectory Bundle Method.
II Roll-to-Roll System Dynamics
Consider a R2R manufacturing line with $N$ motorized rollers indexed by $i\in\{1,\ldots,N\}$, with continuous web material spanning between consecutive rollers. We assume: (1) passive rollers are omitted; (2) web tension $T_{i}$ is spatially uniform between adjacent rollers; (3) no slippage occurs, giving $v_{i}=\omega_{i}R_{i}$; and (4) the web exhibits linear elastic behavior with Young’s modulus $E$ and cross-sectional area $A$.
The tension dynamics in section $i$ follow from mass conservation:
$\frac{\mathrm{d}T_{i}}{\mathrm{d}t}=\frac{EA}{L_{i}}(v_{i}-v_{i-1})+\frac{1}{L_{i}}(T_{i-1}v_{i-1}-T_{i}v_{i}),$ (1)
where $L_{i}$ is the span length and boundary conditions $T_{0}=T_{N+1}=0$ apply. The roller velocity dynamics result from torque balance:
$\frac{\mathrm{d}v_{i}}{\mathrm{d}t}=\frac{R_{i}^{2}}{J_{i}}(T_{i+1}-T_{i})-\frac{f_{i}}{J_{i}}v_{i}+\frac{R_{i}}{J_{i}}u_{i}+b_{i}\frac{\mathrm{d}w_{i}}{\mathrm{d}t},$ (2)
where $J_{i}$ is roller inertia, $f_{i}$ is viscous friction, $u_{i}$ is motor torque, and $w_{i}$ is a Brownian motion capturing stochastic disturbances.
For trajectory optimization, we discretize (1)–(2) using Euler-Maruyama integration with sampling period $\Delta t$. Define state $x_{t}=[T_{1},\ldots,T_{N},v_{1},\ldots,v_{N}]^{\top}\in\mathbb{R}^{2N}$ and control $u_{t}=[u_{1},\ldots,u_{N}]^{\top}\in\mathbb{R}^{N}$. The deterministic propagation map used for bundle construction is:
$F(x_{t},u_{t})\triangleq x_{t}+\big(f(x_{t})+Gu_{t}\big)\Delta t,$ (3)
where $f:\mathbb{R}^{2N}\to\mathbb{R}^{2N}$ encodes the drift terms from (1)–(2) and $G\in\mathbb{R}^{2N\times N}$ maps control inputs to velocity states.
III Adaptive Trajectory Bundle Method: Theoretical Framework
This section develops the adaptive trajectory bundle method for R2R manufacturing control. We first review derivative-free function approximation via trajectory bundles, then formulate the convex subproblem, and finally introduce adaptive mechanisms for trust region and penalty management that eliminate manual tuning.
III-A Notation and Preliminaries
Let $N$ denote the number of motorized rollers and $H$ the prediction horizon. The state dimension is $n_{x}=2N$ (tensions and velocities) and control dimension is $n_{u}=N$ (motor torques). A trajectory over horizon $H$ is denoted $z=\{(x_{k},u_{k})\}_{k=1}^{H}$ where $x_{k}\in\mathbb{R}^{n_{x}}$ and $u_{k}\in\mathbb{R}^{n_{u}}$. We endow the trajectory space with the norm:
$\|z\|=\sqrt{\sum_{k=1}^{H}\left(\|x_{k}\|^{2}+\|u_{k}\|^{2}\right)}.$ (4)
The probability simplex is denoted $\Delta^{m-1}=\{\alpha\in\mathbb{R}^{m}:\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\}$, and $B(z,\Delta)=\{z^{\prime}:\|z^{\prime}-z\|\leq\Delta\}$ is the closed ball of radius $\Delta$ centered at $z$. The operator $\cdot]_{-}:\mathbb{R}^{n}\to\mathbb{R}^{n}_{\geq 0}$ is defined by $v]_{-}=\max(0,-v)$ applied element-wise, extracting the negative part so that $c]_{-}>0$ indicates constraint violation when $c\geq 0$ is required.
III-B Derivative-Free Function Approximation via Bundles
The trajectory bundle method approximates nonlinear functions through interpolation of sampled values rather than Taylor series expansion. This approach requires only function evaluations, making it suitable for systems where analytical gradients are expensive or unavailable [conn2009derivative]. The bundle method terminology originates from nonsmooth convex optimization [lemarechal1975, kiwiel1985], where function values and subgradients are collected into “bundles” to construct piecewise-linear approximations. The trajectory bundle method [tracy2025trajectory] adapts this philosophy to trajectory optimization using function value interpolation rather than subgradient cutting planes.
III-B1 Bundle Interpolation
Consider an arbitrary function $p:\mathbb{R}^{a}\to\mathbb{R}^{b}$. Given a reference point $\bar{y}\in\mathbb{R}^{a}$ and a sampling radius $\Delta>0$, we collect $m$ evaluation points $\{y_{1},\ldots,y_{m}\}\subset B(\bar{y},\Delta)$ and compute $p_{i}=p(y_{i})$ for each sample. These are arranged into matrices:
$W_{y}=\begin{bmatrix}y_{1}&\cdots&y_{m}\end{bmatrix}\in\mathbb{R}^{a\times m},\quad W_{p}=\begin{bmatrix}p_{1}&\cdots&p_{m}\end{bmatrix}\in\mathbb{R}^{b\times m}.$ (5)
The bundle approximation represents candidate solutions as convex combinations of samples [conn2009derivative]:
$y=W_{y}\alpha,\quad p(y)\approx W_{p}\alpha,\quad\alpha\in\Delta^{m-1}.$ (6)
This construction is exact when $p$ is affine. Importantly, restricting $\alpha$ to the simplex automatically confines solutions to $\mathrm{conv}(\{y_{1},\ldots,y_{m}\})\subseteq B(\bar{y},\Delta)$, providing an implicit trust region.
III-B2 Multiple-Shooting Trajectory Representation
For trajectory optimization over horizon $H$, we employ a multiple-shooting discretization [bock1984, betts2010], treating states at discrete time points as independent variables linked through dynamics constraints. At each time index $k\in\{1,\ldots,H\}$, we sample $m$ candidate state-control pairs around the current iterate $(\bar{x}_{k},\bar{u}_{k})$ and evaluate:
$W_{x}^{(k)}$ $=\begin{bmatrix}x_{1}^{(k)}&\cdots&x_{m}^{(k)}\end{bmatrix}\in\mathbb{R}^{n_{x}\times m},$ (7)
$W_{u}^{(k)}$ $=\begin{bmatrix}u_{1}^{(k)}&\cdots&u_{m}^{(k)}\end{bmatrix}\in\mathbb{R}^{n_{u}\times m},$ (8)
$W_{f}^{(k)}$ $=\begin{bmatrix}F(x_{1}^{(k)},u_{1}^{(k)})&\cdots&F(x_{m}^{(k)},u_{m}^{(k)})\end{bmatrix}\in\mathbb{R}^{n_{x}\times m},$ (9)
$W_{r}^{(k)}$ $=\begin{bmatrix}r_{k}(x_{1}^{(k)},u_{1}^{(k)})&\cdots&r_{k}(x_{m}^{(k)},u_{m}^{(k)})\end{bmatrix}\in\mathbb{R}^{n_{r}\times m},$ (10)
$W_{c}^{(k)}$ $=\begin{bmatrix}c_{k}(x_{1}^{(k)},u_{1}^{(k)})&\cdots&c_{k}(x_{m}^{(k)},u_{m}^{(k)})\end{bmatrix}\in\mathbb{R}^{n_{c}\times m},$ (11)
where $F$ is the discrete-time dynamics map (3), $r_{k}$ is the stage cost residual, and $c_{k}$ captures operational constraints at step $k$. Here $c_{k}=(c_{\mathrm{hard},k},c_{1,k},\ldots,c_{J,k})$ decomposes into hard constraints (state/input bounds) and $J$ classes of soft constraints. We denote the corresponding bundle matrices as $W_{c,\mathrm{hard}}^{(k)}$ and $W_{c,j}^{(k)}$ for $j=1,\ldots,J$.
III-C Constrained Bundle Subproblem
Using the bundle matrices, we formulate a convex approximation of the nonlinear trajectory optimization problem. At iteration $\ell$, the subproblem is:
$\operatorname*{minimize}_{\alpha,s,w,d}$ $\sum_{k=1}^{H}\|W_{r}^{(k)}\alpha^{(k)}\|_{2}^{2}+\mu_{\ell}\sum_{k=1}^{H-1}\left(\|s_{k}\|_{1}+\|w_{k}\|_{1}\right)$ (12)
$+\sum_{j=1}^{J}\gamma_{j,\ell}\sum_{k=1}^{H-1}\|d_{k,j}\|_{1}$
$\operatorname{subject~to}$ $W_{f}^{(k)}\alpha^{(k)}=W_{x}^{(k+1)}\alpha^{(k+1)}+s_{k},$
$W_{c,\mathrm{h}}^{(k)}\alpha^{(k)}+w_{k}\geq 0,$
$W_{c,j}^{(k)}\alpha^{(k)}+d_{k,j}\geq 0,\quad j=1,\ldots,J,$
$\alpha^{(k)}\in\Delta^{m-1},\quad k=1,\ldots,H,$
$s_{k}\in\mathbb{R}^{n_{x}},\;w_{k}\geq 0,\;d_{k,j}\geq 0,$
where $s_{k}\in\mathbb{R}^{n_{x}}$ are slack variables for dynamics defects (unconstrained in sign), $w_{k}\in\mathbb{R}_{\geq 0}^{n_{\mathrm{hard}}}$ are non-negative slack variables for hard inequality constraints (state/input bounds), and $d_{k,j}\in\mathbb{R}_{\geq 0}^{n_{j}}$ are non-negative slack variables for soft constraint class $j\in\{1,\ldots,J\}$. The penalty parameter $\mu_{\ell}>0$ penalizes dynamics and hard constraint violations, while $\gamma_{j,\ell}>0$ penalizes soft constraint violations in class $j$. The use of $\ell_{1}$ norms for exact penalty functions follows classical approaches in constrained optimization [han1977, powell1978].
Problem (12) is a convex optimization problem: the objective combines a convex quadratic term with convex $\ell_{1}$ penalties, all constraints are linear, and the feasible region is a Cartesian product of simplices and non-negative orthants.
After solving (12), the candidate trajectory is recovered via:
$x_{k}^{\ell+1}=W_{x}^{(k)}\alpha^{(k)*},\quad u_{k}^{\ell+1}=W_{u}^{(k)}\alpha^{(k)*},\quad k=1,\ldots,H.$ (13)
Remark 1 (Relationship to Original TBM).
The formulation (12) extends the original trajectory bundle method [tracy2025trajectory] in two key aspects:
(i) Multiple Penalty Classes: The original TBM uses a single penalty parameter $\mu$ for all constraints. We introduce constraint classes indexed by $j\in\{1,\ldots,J\}$ with separate penalties $\gamma_{j}$, enabling asymmetric treatment of different violation types. For R2R systems, this is critical for safety: over-tension (web breakage risk) is penalized more heavily ($\gamma_{+}=100$) than under-tension (wrinkling, $\gamma_{-}=10$).
(ii) Adaptive Penalty Scheduling: While the original TBM uses fixed penalties, we allow $\mu_{\ell}$ and $\gamma_{j,\ell}$ to adapt based on observed constraint violations, eliminating the need for manual penalty tuning (see Section III-D below).
III-D Adaptive Mechanisms
The original TBM [tracy2025trajectory] uses fixed penalty parameters throughout optimization. While effective, this requires manual tuning and lacks convergence guarantees. We introduce adaptive mechanisms for trust region and penalty management that automatically adjust based on constraint satisfaction metrics. Classical trust region methods adapt $\Delta$ based on the ratio of actual to predicted objective reduction [conn2000trust]. Our adaptation rules instead use constraint violation metrics, following the philosophy of [tracy2025trajectory] while introducing separate penalty classes for asymmetric constraint treatment.
III-D1 Constraint Violation Metrics
We define separate violation metrics for each constraint type:
Definition 1 (Constraint Violation Metrics).
Given optimal slack variables $(s^{*},w^{*},\{d_{j}^{*}\})$ from subproblem (12) at iteration $\ell$:
$\nu_{\mathrm{dyn}}^{\ell}$ $=\max_{k=1,\ldots,H-1}\|s_{k}^{*}\|_{\infty},$ (14)
$\nu_{\mathrm{hard}}^{\ell}$ $=\max_{k=1,\ldots,H-1}\|w_{k}^{*}\|_{\infty},$ (15)
$\nu_{j}^{\ell}$ $=\sum_{k=1}^{H-1}\|d_{k,j}^{*}\|_{1},\quad j=1,\ldots,J.$ (16)
III-D2 Trust Region Adaptation
The trust region radius $\Delta^{\ell}>0$ controls the locality of the bundle approximation. We adapt it based on constraint satisfaction:
$\Delta^{\ell+1}=\begin{cases}\min(\beta_{\exp}\Delta^{\ell},\Delta_{\max})&\text{if }\nu_{\mathrm{dyn}}^{\ell}<\tau_{\mathrm{feas}}\text{ and }\nu_{\mathrm{hard}}^{\ell}<\tau_{\mathrm{feas}},\\ \max(\beta_{\mathrm{con}}\Delta^{\ell},\Delta_{\min})&\text{if }\nu_{\mathrm{dyn}}^{\ell}>\tau_{\mathrm{viol}}\text{ or }\nu_{\mathrm{hard}}^{\ell}>\tau_{\mathrm{viol}},\\ \Delta^{\ell}&\text{otherwise},\end{cases}$ (17)
where $\beta_{\exp}>1$ is the expansion factor, $\beta_{\mathrm{con}}<1$ is the contraction factor, and $\Delta_{\min},\Delta_{\max}$ bound the trust region radius. The thresholds satisfy $0<\epsilon_{\mathrm{feas}}<\tau_{\mathrm{feas}}<\tau_{\mathrm{viol}}$, where $\epsilon_{\mathrm{feas}}$ is the convergence tolerance, $\tau_{\mathrm{feas}}$ triggers trust region expansion, and $\tau_{\mathrm{viol}}$ triggers contraction and penalty increase.
The rationale is as follows: when violations are small (below $\tau_{\mathrm{feas}}$), the bundle approximation is accurate over a larger region, justifying trust region expansion; when violations are large (above $\tau_{\mathrm{viol}}$), the approximation is poor, so reducing $\Delta$ forces samples closer to $z^{\ell}$ for better local accuracy; when violations are moderate (between $\tau_{\mathrm{feas}}$ and $\tau_{\mathrm{viol}}$), the current sampling is appropriate and the trust region radius is maintained.
III-D3 Penalty Adaptation
Penalty parameters increase when violations persist above thresholds. This finite stabilization property is standard for geometrically increasing penalty sequences [nocedal2006numerical]. Since $\mu$ penalizes both dynamics and hard constraint violations in the subproblem (12), we increase $\mu$ when either type of violation exceeds its threshold:
$\mu^{\ell+1}=\begin{cases}\min(\rho_{\mu}\mu^{\ell},\mu_{\max})&\text{if }\nu_{\mathrm{dyn}}^{\ell}>\tau_{\mathrm{viol}}\text{ or }\nu_{\mathrm{hard}}^{\ell}>\tau_{\mathrm{viol}},\\ \mu^{\ell}&\text{otherwise}.\end{cases}$ (18)
Similarly, soft constraint penalties increase when their respective violations exceed thresholds:
$\gamma_{j}^{\ell+1}=\begin{cases}\min(\rho_{\gamma}\gamma_{j}^{\ell},\gamma_{j,\max})&\text{if }\nu_{j}^{\ell}>\tau_{j},\\ \gamma_{j}^{\ell}&\text{otherwise},\end{cases}\quad j=1,\ldots,J,$ (19)
where $\rho_{\mu},\rho_{\gamma}>1$ are increase factors, and $\tau_{\mathrm{viol}},\tau_{j}$ are violation thresholds.
Algorithm 1 provides the complete specification.
Algorithm 1 Adaptive Trajectory Bundle Method
1:Initial trajectory $z^{0}=\{(x_{k}^{0},u_{k}^{0})\}_{k=1}^{H}$, parameters $(\Delta^{0},\mu^{0},\boldsymbol{\gamma}^{0})$
2:Converged trajectory $z^{*}$
[…]
4:while not converged do
5:  Sample: Generate $\mathcal{Y}^{\ell}=\{y_{1},\ldots,y_{m}\}\subset B(z^{\ell},\Delta^{\ell})$ with $z^{\ell}\in\mathcal{Y}^{\ell}$
6:  Build Bundles: Evaluate $W_{x}^{(k)},W_{u}^{(k)},W_{f}^{(k)},W_{r}^{(k)},W_{c}^{(k)}$ for $k=1,\ldots,H$
7:  Solve Subproblem: Compute $(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})$ from (12)
8:  Recover Trajectory: $z^{\ell+1}\leftarrow\{(W_{x}^{(k)}\alpha^{(k)*},W_{u}^{(k)}\alpha^{(k)*})\}_{k=1}^{H}$
9:  Compute Violations: $\nu_{\mathrm{dyn}}^{\ell}\leftarrow\max_{k}\|s_{k}^{*}\|_{\infty}$, $\nu_{\mathrm{hard}}^{\ell}\leftarrow\max_{k}\|w_{k}^{*}\|_{\infty}$, $\nu_{j}^{\ell}\leftarrow\sum_{k}\|d_{k,j}^{*}\|_{1}$
10:  Adapt Trust Region: Update $\Delta^{\ell+1}$ via (17)
11:  Adapt Penalties: Update $\mu^{\ell+1}$ via (18) and $\gamma_{j}^{\ell+1}$ via (19)
12:  Check Convergence: If $\nu_{\mathrm{dyn}}^{\ell}<\epsilon_{\mathrm{feas}}$ and $\nu_{\mathrm{hard}}^{\ell}<\epsilon_{\mathrm{feas}}$ and $\|z^{\ell+1}-z^{\ell}\|<\epsilon_{z}$: break
13:  $\ell\leftarrow\ell+1$
14:end while
15:return $z^{\ell+1}$
IV Convergence Analysis
This section establishes convergence guarantees for Algorithm 1. We prove that the adaptive mechanisms drive constraint violations to zero and the algorithm converges to stationary points of the penalized objective, which coincide with stationary points of the original constrained R2R control problem under sufficient penalty magnitudes. The proof structure follows classical trust region convergence theory [conn2000trust], adapted to the penalty formulation and always-accept strategy of [tracy2025trajectory].
We first define the penalized objective function that underlies the convergence analysis.
Definition 2 (Penalized Objective).
For trajectory $z=\{(x_{k},u_{k})\}_{k=1}^{H}$ and penalty parameters $(\mu,\boldsymbol{\gamma})=(\mu,\gamma_{1},\ldots,\gamma_{J})$, define:
$\phi_{\mu,\boldsymbol{\gamma}}(z)=$ $\sum_{k=1}^{H}\|r_{k}(x_{k},u_{k})\|_{2}^{2}$ (20)
$+\mu\sum_{k=1}^{H-1}\|x_{k+1}-F(x_{k},u_{k})\|_{1}$
$+\mu\sum_{k=1}^{H-1}\|[c_{\mathrm{hard},k}(x_{k},u_{k})]_{-}\|_{1}$
$+\sum_{j=1}^{J}\gamma_{j}\sum_{k=1}^{H-1}\|[c_{j,k}(x_{k},u_{k})]_{-}\|_{1},$
where $v]_{-}=\max(0,-v)$ is applied element-wise, extracting constraint violations.
Assumption 1 (Lipschitz Continuity).
The stage cost residual $r_{k}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{r}}$, dynamics map $F:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{x}}$, and constraint functions $c_{k}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{c}}$ are Lipschitz continuous with constants $L_{r}$, $L_{F}$, and $L_{c}$ respectively.
Assumption 2 (Bounded Level Sets).
For any fixed penalties $(\mu,\boldsymbol{\gamma})$ with $\mu\leq\mu_{\max}$ and $\gamma_{j}\leq\gamma_{j,\max}$, the level set
$\mathcal{L}(\mu,\boldsymbol{\gamma})=\left\{z:\phi_{\mu,\boldsymbol{\gamma}}(z)\leq\phi_{\mu,\boldsymbol{\gamma}}(z^{0})\right\}$ (21)
is compact. Furthermore, the cost residuals are bounded on this level set: $\|r_{k}(x_{k},u_{k})\|\leq R$ for all $z\in\mathcal{L}(\mu,\boldsymbol{\gamma})$ and all $k$.
Assumption 3 (Controlled Sampling).
At each iteration $\ell$, the sampling procedure generates $\mathcal{Y}^{\ell}=\{y_{1},\ldots,y_{m}\}$ satisfying: (i) $y_{i}\in B(z^{\ell},\Delta^{\ell})$ for all $i\in\{1,\ldots,m\}$; (ii) $z^{\ell}\in\mathcal{Y}^{\ell}$, i.e., the current iterate is included; and (iii) when $\Delta^{\ell}>0$, samples provide sufficient directional coverage to detect descent directions.
Assumption 4 (Sufficient Trust Region Contraction).
The minimum trust region radius $\Delta_{\min}$ is chosen small enough such that for any trajectory $z$ in the level set $\mathcal{L}(\mu_{\max},\boldsymbol{\gamma}_{\max})$, if the bundle approximation is constructed with radius $\Delta_{\min}$, then the approximation error satisfies $CL_{\phi}\Delta_{\min}<\epsilon_{\mathrm{feas}}$, where $C$ is the constant from Lemma 3 and $\epsilon_{\mathrm{feas}}$ is the convergence tolerance.
Definition 3 (Local Stationarity).
A trajectory $z^{*}$ is stationary for $\phi_{\mu,\boldsymbol{\gamma}}$ if for all sufficiently small $\Delta>0$ and any sample set $\mathcal{Y}\subset B(z^{*},\Delta)$ with $z^{*}\in\mathcal{Y}$, solving the subproblem (12) yields:
$\lim_{\Delta\to 0}\|z^{+}-z^{*}\|=0,$ (22)
where $z^{+}=W_{z}\alpha^{*}$ is the recovered trajectory.
Remark 2.
This definition captures the practical notion that no local convex combination of nearby samples improves the objective. It is the appropriate stationarity concept for derivative-free optimization [conn2009derivative].
Lemma 1 (Lipschitz Constant).
Under Assumptions 1 and 2, the penalized objective $\phi_{\mu,\boldsymbol{\gamma}}$ is Lipschitz continuous on $\mathcal{L}(\mu,\boldsymbol{\gamma})$ with constant:
$L_{\phi}(\mu,\boldsymbol{\gamma})=2HRL_{r}+(H-1)\mu(1+L_{F}+L_{c})+(H-1)L_{c}\sum_{j=1}^{J}\gamma_{j},$ (23)
where $R$ is the bound on cost residuals from Assumption 2.
Proof.
The proof follows from standard Lipschitz composition rules; see, e.g., [conn2009derivative, Chapter 1]. The $\ell_{1}$ norm and $\cdot$$_{-}$ operator are Lipschitz with constant 1. For the cost term, using $|\|a\|^{2}-\|b\|^{2}|=|(\|a\|+\|b\|)|\cdot|\|a\|-\|b\||\leq(\|a\|+\|b\|)\|a-b\|$ and the bound $\|r_{k}\|\leq R$:
$\left|\|r_{k}(x_{k},u_{k})\|^{2}-\|r_{k}(x_{k}^{\prime},u_{k}^{\prime})\|^{2}\right|$ $\leq 2R\|r_{k}(x_{k},u_{k})-r_{k}(x_{k}^{\prime},u_{k}^{\prime})\|$ (24)
$\leq 2RL_{r}\|(x_{k},u_{k})-(x_{k}^{\prime},u_{k}^{\prime})\|.$
Summing over $k=1,\ldots,H$ gives the first term $2HRL_{r}$.
For the dynamics penalty term $\mu\|x_{k+1}-F(x_{k},u_{k})\|_{1}$:
$\begin{split}&\left|\|x_{k+1}-F(x_{k},u_{k})\|_{1}-\|x^{\prime}_{k+1}-F(x^{\prime}_{k},u^{\prime}_{k})\|_{1}\right|\\ &\quad\leq\|(x_{k+1}-x^{\prime}_{k+1})-(F(x_{k},u_{k})-F(x^{\prime}_{k},u^{\prime}_{k}))\|_{1}\\ &\quad\leq\|x_{k+1}-x^{\prime}_{k+1}\|_{1}+L_{F}\|(x_{k},u_{k})-(x^{\prime}_{k},u^{\prime}_{k})\|.\end{split}$ (25)
This contributes $(1+L_{F})$ per time step. Similar analysis for constraint terms yields the stated bound. ∎
Lemma 2 (Penalty Stabilization).
Under Algorithm 1, there exists $K^{*}<\infty$ such that $\mu^{\ell}=\bar{\mu}$ and $\gamma_{j}^{\ell}=\bar{\gamma}_{j}$ for all $\ell\geq K^{*}$, where:
$K^{*}\leq\left\lceil\log_{\rho_{\mu}}\left(\frac{\mu_{\max}}{\mu^{0}}\right)\right\rceil+\sum_{j=1}^{J}\left\lceil\log_{\rho_{\gamma}}\left(\frac{\gamma_{j,\max}}{\gamma_{j}^{0}}\right)\right\rceil.$ (26)
Proof.
Penalties increase geometrically by factors $\rho_{\mu}$ or $\rho_{\gamma}>1$ in the Adapt Penalties step when violations exceed thresholds, and are capped at $\mu_{\max},\gamma_{j,\max}$. Each penalty reaches its maximum after at most $\lceil\log_{\rho}(\text{max}/\text{init})\rceil$ increases. This finite stabilization property is standard for geometrically increasing penalty sequences; cf. [nocedal2006numerical, Section 17.2]. ∎
Lemma 3 (Bundle Approximation Accuracy).
Let $z^{+}=W_{z}\alpha^{*}$ be the recovered trajectory from subproblem (12). Under Assumptions 13, the subproblem objective approximates the penalized objective with error:
$\left|\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{+})-J_{\mathrm{sub}}(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})\right|\leq CL_{\phi}\Delta^{\ell},$ (27)
for some constant $C>0$ depending only on problem dimensions, where $J_{\mathrm{sub}}$ denotes the subproblem objective.
Proof.
This bound follows from standard interpolation error analysis in derivative-free optimization [conn2009derivative, Chapter 10]. The subproblem approximates the cost via $\sum_{k}\|W_{r}^{(k)}\alpha^{(k)}\|^{2}\approx\sum_{k}\|r_{k}(x_{k}^{+},u_{k}^{+})\|^{2}$ with interpolation error $O(\Delta^{\ell})$ by Lipschitz continuity. Slack variables bound exact violations: $\|x_{k+1}^{+}-F(x_{k}^{+},u_{k}^{+})\|\leq\|s_{k}^{*}\|+O(\Delta^{\ell})$ since $W_{f}^{(k)}\alpha^{(k)}$ approximates $F(x_{k}^{+},u_{k}^{+})$ with error $O(\Delta^{\ell})$. Combining these approximations yields the stated bound. ∎
Lemma 4 (Descent Property).
At each iteration $\ell$, the algorithm satisfies:
$\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{\ell+1})\leq\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{\ell})+CL_{\phi}\Delta^{\ell}.$ (28)
Proof.
The bounded increase property follows from the subproblem including the current iterate as a feasible point, a standard technique in trust region methods [conn2000trust]. Since $z^{\ell}\in\mathcal{Y}^{\ell}$ (Assumption 3(ii)), we can represent it with weights $\alpha_{i}=1$ for the index where $y_{i}=z^{\ell}$ and $\alpha_{j}=0$ otherwise. For this choice of $\alpha$, the dynamics constraint in (12) requires slack $s_{k}=F(x_{k}^{\ell},u_{k}^{\ell})-x_{k+1}^{\ell}$, which equals the dynamics defect of the current iterate. Similarly, $w_{k}$ and $d_{k,j}$ equal the constraint violations of $z^{\ell}$. Thus, the subproblem objective at this feasible point equals:
$J_{\mathrm{sub}}(\alpha^{\ell},s^{\ell},w^{\ell},\{d_{j}^{\ell}\})=\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{\ell})+O(\Delta^{\ell}),$ (29)
where the $O(\Delta^{\ell})$ term accounts for bundle interpolation error (Lemma 3).
Since the subproblem minimizes over all $\alpha\in\Delta^{m-1}$ and associated slacks:
$\begin{split}J_{\mathrm{sub}}(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})&\leq J_{\mathrm{sub}}(\alpha^{\ell},s^{\ell},w^{\ell},\{d_{j}^{\ell}\})\\ &\leq\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{\ell})+O(\Delta^{\ell}).\end{split}$ (30)
By Lemma 3, $\phi_{\mu^{\ell},\boldsymbol{\gamma}^{\ell}}(z^{\ell+1})\leq J_{\mathrm{sub}}(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})+CL_{\phi}\Delta^{\ell}$, which yields the result. ∎
Lemma 5 (Feasibility Improvement).
Suppose $\ell\geq K^{*}$ (penalties have stabilized at $\bar{\mu},\bar{\boldsymbol{\gamma}}$) and either $\nu_{\mathrm{dyn}}^{\ell}>\tau_{\mathrm{viol}}$ or $\nu_{\mathrm{hard}}^{\ell}>\tau_{\mathrm{viol}}$. Then the penalized objective satisfies:
$\phi_{\bar{\mu},\bar{\boldsymbol{\gamma}}}(z^{\ell+1})<\phi_{\bar{\mu},\bar{\boldsymbol{\gamma}}}(z^{\ell}),$ (31)
provided the trust region $\Delta^{\ell}$ is small enough that the approximation error $CL_{\phi}\Delta^{\ell}<\bar{\mu}\tau_{\mathrm{viol}}$.
Proof.
This feasibility improvement property is classical for exact penalty methods; see [nocedal2006numerical, Theorem 17.3]. When violations exceed $\tau_{\mathrm{viol}}$, the trust region contracts via (17), forcing samples closer to the current iterate. The subproblem finds a trajectory $z^{\ell+1}$ that reduces violations while maintaining or improving cost, since the $\ell_{1}$ penalty terms in (12) drive the optimizer to reduce $\|s_{k}\|_{1}$ and $\|w_{k}\|_{1}$.
Since the current iterate $z^{\ell}$ is feasible for the subproblem (Lemma 4), and the subproblem minimizes, we have:
$J_{\mathrm{sub}}(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})\leq J_{\mathrm{sub}}(\alpha^{\ell},s^{\ell},w^{\ell},\{d_{j}^{\ell}\}).$ (32)
When $\nu_{\mathrm{dyn}}^{\ell}>\tau_{\mathrm{viol}}$ or $\nu_{\mathrm{hard}}^{\ell}>\tau_{\mathrm{viol}}$, the penalty contribution from violations is at least $\bar{\mu}\tau_{\mathrm{viol}}$. The subproblem actively minimizes this term, guaranteeing a reduction of at least $\bar{\mu}\tau_{\mathrm{viol}}-CL_{\phi}\Delta^{\ell}>0$ in the true penalized objective when $CL_{\phi}\Delta^{\ell}<\bar{\mu}\tau_{\mathrm{viol}}$. ∎
Lemma 6 (Trust Region Contraction Effect).
Under Assumption 4, when $\Delta^{\ell}=\Delta_{\min}$, the subproblem slacks satisfy:
$\|s_{k}^{*}\|_{\infty}<\tau_{\mathrm{feas}},\quad\|w_{k}^{*}\|_{\infty}<\tau_{\mathrm{feas}},\quad\text{for all }k,$ (33)
provided $z^{\ell}$ is in the level set $\mathcal{L}(\mu_{\max},\boldsymbol{\gamma}_{\max})$.
Proof.
The argument follows the classical trust region strategy that sufficiently small trust regions yield accurate models [conn2000trust, Chapter 6]. When $\Delta^{\ell}=\Delta_{\min}$, all samples including $z^{\ell}$ lie within a ball of radius $\Delta_{\min}$ where the bundle approximation has error $O(\Delta_{\min})$. The subproblem’s $\ell_{1}$ penalty terms actively minimize slack variables. Since the optimizer can choose any convex combination of samples, and by Assumption 4 the approximation error satisfies $CL_{\phi}\Delta_{\min}<\epsilon_{\mathrm{feas}}<\tau_{\mathrm{feas}}$, the optimal slacks satisfy $\|s_{k}^{*}\|_{\infty}\leq CL_{\phi}\Delta_{\min}<\tau_{\mathrm{feas}}$ and similarly $\|w_{k}^{*}\|_{\infty}<\tau_{\mathrm{feas}}$. ∎
Remark 3 (Violation Metric Design).
Definition 1 uses different norms for different constraint types: $\nu_{\mathrm{dyn}}^{\ell}$ and $\nu_{\mathrm{hard}}^{\ell}$ take the maximum over time of the $\|\cdot\|_{\infty}$ norm (worst-case violation at any time step), while $\nu_{j}^{\ell}$ uses $\sum_{k}\|\cdot\|_{1}$ (cumulative violation). This asymmetry reflects the design philosophy that dynamics and hard constraints (state/input bounds) must be satisfied uniformly at every time step, whereas soft constraints may tolerate occasional small violations provided the cumulative violation remains bounded. Consequently, thresholds $\tau_{\mathrm{viol}}$ and $\tau_{j}$ should be calibrated with this distinction in mind.
Theorem 1 (Convergence to Feasible Stationary Points).
Under Assumptions 14, suppose the penalty bounds satisfy $\mu_{\max}$ and $\gamma_{j,\max}$ large enough such that the exact penalty equivalence holds (i.e., constrained stationary points are also stationary for the penalized problem when $\mu\geq\mu_{\max}$ and $\gamma_{j}\geq\gamma_{j,\max}$). Then Algorithm 1 satisfies:
  1. (i)
    Penalty Stabilization: There exists $K^{*}<\infty$ such that $(\mu^{\ell},\boldsymbol{\gamma}^{\ell})=(\bar{\mu},\bar{\boldsymbol{\gamma}})$ for all $\ell\geq K^{*}$.
  2. (ii)
    Feasibility Attainment: $\lim_{\ell\to\infty}(\nu_{\mathrm{dyn}}^{\ell}+\nu_{\mathrm{hard}}^{\ell})=0$.
  3. (iii)
    Trajectory Convergence: $\lim_{\ell\to\infty}\|z^{\ell+1}-z^{\ell}\|=0$.
  4. (iv)
    Stationarity: Every accumulation point of $\{z^{\ell}\}$ is stationary for $\phi_{\bar{\mu},\bar{\boldsymbol{\gamma}}}$ in the sense of Definition 3, and hence is a stationary point of the original constrained R2R control problem.
Proof.
Part (i): Immediate from Lemma 2.
Part (ii): For $\ell\geq K^{*}$, denote $\bar{\phi}=\phi_{\bar{\mu},\bar{\boldsymbol{\gamma}}}$. Suppose violations remain above $\tau_{\mathrm{viol}}$ infinitely often. Each time $\nu_{\mathrm{dyn}}^{\ell}>\tau_{\mathrm{viol}}$ or $\nu_{\mathrm{hard}}^{\ell}>\tau_{\mathrm{viol}}$, the trust region contracts via (17). After at most $\lceil\log_{\beta_{\mathrm{con}}}(\Delta_{\min}/\Delta^{0})\rceil$ contractions, $\Delta^{\ell}=\Delta_{\min}$.
Once $\Delta^{\ell}=\Delta_{\min}$, Lemma 6 (using Assumption 4) ensures that subproblem slacks satisfy $\|s_{k}^{*}\|_{\infty}<\tau_{\mathrm{feas}}<\tau_{\mathrm{viol}}$ and $\|w_{k}^{*}\|_{\infty}<\tau_{\mathrm{feas}}<\tau_{\mathrm{viol}}$. This contradicts the assumption that violations remain above $\tau_{\mathrm{viol}}$.
Therefore, after finitely many iterations, violations drop below $\tau_{\mathrm{viol}}$ and remain there. Once violations are below $\tau_{\mathrm{viol}}$, Lemma 4 ensures the penalized objective $\bar{\phi}$ is non-increasing (up to $O(\Delta^{\ell})$ error). Since $\bar{\phi}$ includes penalty terms $\bar{\mu}\sum_{k}(\|s_{k}\|_{1}+\|w_{k}\|_{1})$, and $\bar{\phi}$ is bounded below on the compact level set (Assumption 2), the sequence $\sum_{k}(\|s_{k}\|_{1}+\|w_{k}\|_{1})$ must converge to zero, implying $\nu_{\mathrm{dyn}}^{\ell}\to 0$ and $\nu_{\mathrm{hard}}^{\ell}\to 0$.
Part (iii): We show trajectory convergence by contradiction. Let $\bar{\phi}^{*}=\liminf_{\ell\to\infty}\bar{\phi}(z^{\ell})$, which exists since $\bar{\phi}$ is bounded below on the compact level set (Assumption 2).
Suppose $\|z^{\ell+1}-z^{\ell}\|\not\to 0$. Then there exists $\delta>0$ and a subsequence $\{\ell_{k}\}$ such that $\|z^{\ell_{k}+1}-z^{\ell_{k}}\|\geq\delta$ for all $k$.
At each such iteration, the subproblem found a solution $z^{\ell_{k}+1}$ significantly different from $z^{\ell_{k}}$. By Assumption 3(iii), if a direction of improvement exists within radius $\Delta^{\ell_{k}}$, the sampling provides sufficient coverage to detect it. The fact that the optimizer chose $z^{\ell_{k}+1}$ with $\|z^{\ell_{k}+1}-z^{\ell_{k}}\|\geq\delta$ implies a reduction in the subproblem objective. This reduction can come from either the quadratic cost term or the penalty terms (constraint violations).
Once violations are below $\tau_{\mathrm{feas}}$ (Part (ii)), the trust region stabilizes or expands. With bounded $\Delta^{\ell}\leq\Delta_{\max}$, the approximation errors are uniformly bounded: $CL_{\phi}\Delta^{\ell}\leq CL_{\phi}\Delta_{\max}$.
For the penalized objective, combining Lemma 3 with the subproblem optimality:
$\begin{split}\bar{\phi}(z^{\ell_{k}+1})&\leq J_{\mathrm{sub}}(\alpha^{*},s^{*},w^{*},\{d_{j}^{*}\})+CL_{\phi}\Delta^{\ell_{k}}\\ &\leq\bar{\phi}(z^{\ell_{k}})+2CL_{\phi}\Delta_{\max}.\end{split}$ (34)
Since $\bar{\phi}$ is bounded below and the sequence $\{\bar{\phi}(z^{\ell})\}$ cannot decrease by more than $O(\Delta_{\max})$ infinitely often while maintaining $\|z^{\ell+1}-z^{\ell}\|\geq\delta$, we reach a contradiction. The penalized objective would eventually need to decrease below its infimum, which is impossible. Therefore, $\|z^{\ell+1}-z^{\ell}\|\to 0$.
Part (iv): By compactness (Assumption 2), $\{z^{\ell}\}_{\ell\geq K^{*}}$ has accumulation points. Consider any accumulation point $z^{*}$ along a subsequence $z^{\ell_{k}}\to z^{*}$. By Part (iii), $z^{\ell_{k}+1}\to z^{*}$ as well.
For any $\delta>0$, eventually $\|z^{\ell_{k}}-z^{*}\|<\delta/2$. Consider two cases.
Case 1: Trust region bounded away from zero. If $\Delta^{\ell_{k}}\geq\delta^{\prime}>0$ for some $\delta^{\prime}$, then samples $\mathcal{Y}^{\ell_{k}}\subset B(z^{*},\delta)$ provide directional coverage (Assumption 3(iii)). Along the subsequence, $\|z^{\ell_{k}+1}-z^{\ell_{k}}\|\to 0$, meaning the subproblem cannot find improving directions. By Definition 3, $z^{*}$ is stationary for $\bar{\phi}$.
Case 2: Trust region contracted to $\Delta_{\min}$. By Part (ii), $z^{\ell_{k}}$ achieves $\nu_{\mathrm{dyn}}^{\ell_{k}}<\epsilon_{\mathrm{feas}}$ and $\nu_{\mathrm{hard}}^{\ell_{k}}<\epsilon_{\mathrm{feas}}$, so it is nearly feasible. By Lemma 6, small perturbations within radius $\Delta_{\min}$ yield no improvement, again establishing stationarity.
In both cases, $z^{*}$ is stationary for $\bar{\phi}$. By the exact penalty assumption, when $\bar{\mu}=\mu_{\max}$ and $\bar{\gamma}_{j}=\gamma_{j,\max}$, stationary points of $\bar{\phi}$ coincide with stationary points of the constrained problem [nocedal2006numerical, Theorem 17.3]. ∎
Remark 4 (Soft Constraint Violations).
Theorem 1 proves convergence of dynamics and hard constraint violations to zero, but makes no claim about soft constraint violations $\nu_{j}^{\ell}$. This is intentional: soft constraints are penalized but may remain violated at the optimal trade-off between cost minimization and constraint satisfaction. Similarly, Algorithm 1 checks convergence only for dynamics and hard constraints. If exact soft constraint satisfaction is required, the corresponding $\gamma_{j,\max}$ should be set large enough to ensure violations are driven to zero, or the constraint should be reclassified as hard.
Remark 5 (Comparison to Classical Results).
Theorem 1 provides convergence guarantees analogous to classical trust region methods [conn2000trust], but achieves them through adaptive penalty growth and trust region management based on constraint violations rather than acceptance ratios. This simplification maintains theoretical guarantees while improving computational efficiency.
Corollary 1 (Finite-Time Feasibility).
Under the conditions of Theorem 1, Algorithm 1 achieves $\nu_{\mathrm{dyn}}^{\ell}<\epsilon_{\mathrm{feas}}$ and $\nu_{\mathrm{hard}}^{\ell}<\epsilon_{\mathrm{feas}}$ after finitely many iterations. Specifically, feasibility is achieved within:
$K_{\mathrm{feas}}\leq K^{*}+K_{\Delta}$ (35)
iterations, where $K^{*}\leq\lceil\log_{\rho_{\mu}}(\mu_{\max}/\mu^{0})\rceil+\sum_{j=1}^{J}\lceil\log_{\rho_{\gamma}}(\gamma_{j,\max}/\gamma_{j}^{0})\rceil$ bounds penalty stabilization and $K_{\Delta}=\lceil\log_{\beta_{\mathrm{con}}}(\Delta_{\min}/\Delta^{0})\rceil$ bounds trust region contractions.
Proof.
Part (i) of Theorem 1 shows penalties stabilize within $K^{*}$ iterations. Part (ii) shows that violations drop below $\tau_{\mathrm{viol}}$ after at most $K_{\Delta}$ additional trust region contractions. Once $\Delta^{\ell}=\Delta_{\min}$, Lemma 6 ensures violations are below $\tau_{\mathrm{feas}}$, and the algorithm terminates when they fall below $\epsilon_{\mathrm{feas}}<\tau_{\mathrm{feas}}$. ∎
V Simulation Results
V-A Simulation Setup
The proposed Adaptive TBM was evaluated on a simulated roll-to-roll system with $N=6$ motorized rollers and coupled web spans, representing a typical industrial functional film processing line.
Reference roller velocities $v_{i}^{r}(t)$ are computed from mass conservation to ensure equilibrium:
$v_{i}^{r}(t)=\frac{EA-T_{i-1}(t)}{EA-T_{i}(t)}v_{i-1}^{r}(t),\quad v_{0}^{r}=v_{0}(t),$ (36)
where $E$ denotes Young’s modulus, $A$ the cross-sectional area, and $T_{i}(t)$ the tension setpoint for span $i$.
V-B Implementation Details
The Adaptive TBM was implemented with prediction horizon $H=15$ steps ($0.15$s). Following the formulation in Section III, at each timestep the controller constructs bundle matrices and solves the convex subproblem (12):
$\min_{\alpha,s,w,d^{\pm}}$ $\sum_{k=1}^{H}\|W_{r}^{(k)}\alpha^{(k)}\|_{2}^{2}+\mu\sum_{k=1}^{H-1}\left(\|s_{k}\|_{1}+\|w_{k}\|_{1}\right)$ (37)
$+\gamma_{+}\sum_{k=1}^{H-1}\mathbf{1}^{\top}d_{k}^{+}+\gamma_{-}\sum_{k=1}^{H-1}\mathbf{1}^{\top}d_{k}^{-}$
$\operatorname{subject~to}$ $W_{f}^{(k)}\alpha^{(k)}=W_{x}^{(k+1)}\alpha^{(k+1)}+s_{k},$
$W_{c,\mathrm{hard}}^{(k)}\alpha^{(k)}+w_{k}\geq 0,$
$W_{c,+}^{(k)}\alpha^{(k)}+d_{k}^{+}\geq 0,$
$W_{c,-}^{(k)}\alpha^{(k)}+d_{k}^{-}\geq 0,$
$\alpha^{(k)}\in\Delta^{m-1},\quad k=1,\ldots,H,$
where the cost residual bundles $W_{r}^{(k)}$ encode the quadratic tracking objectives:
$r_{k}(x_{k},u_{k})=\begin{bmatrix}Q^{1/2}(x_{k}-x_{k}^{r})\\ R^{1/2}(u_{k}-u_{k}^{r})\\ S^{1/2}\Delta u_{k}\end{bmatrix}$. (38)
The hard constraint bundles $W_{c,\mathrm{hard}}^{(k)}$ encode state and input bounds $(x_{\min}\leq x_{k}\leq x_{\max},u_{\min}\leq u_{k}\leq u_{\max})$, while soft constraint bundles $W_{c,+}^{(k)}$ and $W_{c,-}^{(k)}$ capture asymmetric tension bounds $(T_{\mathrm{ref}}-\delta^{-}\leq T(x_{k})\leq T_{\mathrm{ref}}+\delta^{+})$. The trajectory is recovered via (13): $x_{k}^{*}=W_{x}^{(k)}\alpha^{(k)*}$, $u_{k}^{*}=W_{u}^{(k)}\alpha^{(k)*}$.
Asymmetric penalties $\gamma_{+}=100$ and $\gamma_{-}=10$ prioritize over-tension avoidance (breakage risk) over under-tension (wrinkling), following the multiple penalty class extension in Remark 1. State weights $Q=\mathrm{diag}(100I_{N},10I_{N})$ emphasize tension regulation, while rate penalty $S=0.1I$ ensures smooth actuation.
The implementation follows Algorithm 1 with three key components:
Bundle Generation
At each time index $k$, $m=6N+21$ samples are generated via a hybrid strategy satisfying Assumption 3. A deterministic stencil of $6N$ samples perturbs the nominal $(\bar{x}_{k},\bar{u}_{k})$ by $\pm\Delta^{\ell}$ along each coordinate axis, ensuring sufficient directional coverage. An additional 20 stochastic samples drawn from $\mathcal{N}((\bar{x}_{k},\bar{u}_{k}),\Sigma)$ enhance convex hull coverage. The current iterate is always included in the sample set, as required by Assumption 3(ii).
Convex Subproblem
The bundle matrices $(W_{x}^{(k)},W_{u}^{(k)},W_{f}^{(k)},W_{r}^{(k)},W_{c}^{(k)})$ are constructed via (7)–(11), transforming the nonlinear R2R control problem into the convex subproblem (37) over simplex-constrained interpolation weights $\alpha^{(k)}$. As noted in Section III, the subproblem is convex: the objective combines a convex quadratic term with convex $\ell_{1}$ penalties, and all constraints are linear over the Cartesian product of simplices and non-negative orthants. The subproblem is solved using a standard convex optimization solver in 10–15 ms per iteration.
Adaptive Trust Region and Penalties
Following Definition 1, constraint violations are computed as $\nu_{\mathrm{dyn}}^{\ell}=\max_{k}\|s_{k}^{*}\|_{\infty}$ and $\nu_{\mathrm{hard}}^{\ell}=\max_{k}\|w_{k}^{*}\|_{\infty}$. The trust region adapts via (17) with expansion factor $\beta_{\exp}=1.5$, contraction factor $\beta_{\mathrm{con}}=0.5$, feasibility threshold $\tau_{\mathrm{feas}}=10^{-4}$, and violation threshold $\tau_{\mathrm{viol}}=10^{-2}$. Penalties adapt via (18)–(19) with increase factor $\rho_{\mu}=\rho_{\gamma}=2.0$. This enables robust handling of aggressive transients without manual tuning, as guaranteed by Theorem 1.
V-C Tension Step Tracking
This test evaluates tracking of a tension step change while maintaining regulation in coupled zones. Webs 1, 2, 4, 5, and 6 hold constant references of $28$, $36$, $40$, $24$, and $32\,\text{N}$, respectively. Web 3 undergoes a step from $20\,\text{N}$ to $44\,\text{N}$ at $t=0.5\,\text{s}$, with unwinding velocity fixed at $0.01\,\text{m/s}$.
Fig. 2a shows the tension response. The Adaptive TBM tracks the $24\,\text{N}$ step in Web 3 with rise time comparable to MPC while effectively rejecting coupling disturbances in downstream webs ($T_{4}$$T_{6}$), where MPPI-based methods exhibit sustained oscillations. Fig. 2b confirms that TBM generates smooth torque commands, avoiding the high-frequency chatter observed in MPPI responses.
(a) Tension response.
(b) Control inputs.
Figure 2: Step tracking performance comparison.
V-D Velocity Setpoint Change
This test evaluates tension regulation during a velocity transient. All webs maintain reference tension of $30\,\text{N}$ while unwinding velocity increases from $0.01\,\text{m/s}$ to $0.10\,\text{m/s}$ at $t=0.5\,\text{s}$.
Fig. 3a shows that while all controllers experience transient tension drops due to acceleration, Adaptive TBM stabilizes downstream tensions ($T_{4}$$T_{6}$) faster than MPPI-based baselines. Fig. 3b confirms smooth torque generation similar to MPC.
(a) Tension response.
(b) Control inputs.
Figure 3: Velocity change performance comparison.
Table I summarizes the tension RMSE across both test scenarios. Adaptive TBM achieves the lowest RMSE in both cases, outperforming gradient-based MPC by $4.3\%$ in tension tracking and $11.1\%$ in velocity transient regulation, while maintaining smooth actuation without derivative computation.
TABLE I: Tension RMSE Comparison (N)
Method Tension Step Velocity Change
MPC 1.384 0.879
MPPI 1.419 0.868
MPPI+L1 1.372 0.815
TBM 1.387 0.824
Adaptive TBM 1.324 0.781
V-E Convergence Comparison
Figure 4 compares the convergence behavior of TBM and Adaptive TBM. The adaptive variant reduces peak transient error by 7.4% (3.15 N vs 3.4 N) in tension step tracking, as the trust region mechanism constrains step sizes during large disturbances. Additionally, Adaptive TBM converges approximately 50 iterations faster, benefiting from trust region expansion during refinement. Both methods achieve comparable steady-state accuracy (1.4 N and 0.8 N for the two scenarios), demonstrating that the adaptive mechanisms improve transient response without sacrificing final performance.
Figure 4: Convergence comparison between TBM and Adaptive TBM.
VI Conclusion
This paper developed an adaptive trajectory bundle method for R2R manufacturing control. The proposed framework combines the derivative-free simplicity of sampling-based methods with the constraint-handling rigor of sequential convex programming. Adaptive trust region and penalty mechanisms eliminate manual tuning while ensuring robust convergence. Simulation results on a six-zone R2R system demonstrate tracking accuracy comparable to gradient-based MPC and superior disturbance rejection over MPPI baselines. The approach offers a practical solution for industrial web handling where analytical gradients are unavailable or computationally prohibitive.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.
VII Biography Section
[…]
If you include a photo:
[…]
If you will not include a photo:
[…]